OpenStudy (anonymous):

Which expressions show how 5 • 37 can be rewritten using the distributive property? Choose all answers that are correct. A. 5 • 30 + 7 B. 5 • 30 + 5 • 7 C. 5 • 40 – 5 • 7 D. 5 • 40 – 5 • 3

OpenStudy (mathmate):

Examples for \(4\times 13\) are \( 4\times 7 + 4\times 6\) \( 4\times 14 - 4\times 1\) \( 4\times 10 + 4\times 3\) etc.

OpenStudy (anonymous):

thx

OpenStudy (mathmate):

you're welcome! :)